# ArrowGEV: Grounding Events in Video via Learning the Arrow of Time

### v1

@@ 4 Experiments
Benchmarks. We evaluate our model on three GEV benchmarks: Charades-STA (Sigurdsson et al., 2016), ActivityNet (Caba Heilbron et al., 2015), and TVGBench (Wang et al., 2025c). To further evaluate the generalization ability, we further compare ArrowGEV on the video understanding and reasoning benchmarks, including TempCompass (Liu et al., 2024), MVBench (Li et al., 2024b), VSI-Bench (Yang et al., 2025c), Video-MMMU (Hu et al., 2025), MMVU (Zhao et al., 2025), and VideoMME (Fu et al., 2024).
@@
where $\text{R1@m}(\text{fwd})$ denotes R1@m when predictions align with the ground truth $\mathcal{T}_{gt}$ on the forward video $\mathcal{V}$, and $\text{R1@m}(\text{rev})$ denotes R1@m on the reversed video $\mathcal{V}^{\prime}$ with respect to the mirrored ground truth $\mathcal{R}(\mathcal{T}_{gt})$. The interpretation of the TDD depends on the event category. For Time-sensitive events, ideal models should accurately localize the forward event ($\text{R1@m}(\text{fwd})\rightarrow 1$) but recognize its absence upon reversal due to the nonexistence of the event ($\text{R1@m}(\text{rev})\rightarrow 0$), yielding a TDD approaching 1. For Time-insensitive events, models should demonstrate temporal invariance by consistently localizing the event in both directions ($\text{R1@m}(\text{fwd})\approx\text{R1@m}(\text{rev})$), yielding a TDD approaching 0. This demonstrates that the model correctly recognizes the event’s invariance to temporal direction and exhibits strong consistency.
@@
TVG-R1 (Chen et al., 2025a) 60.7 36.1 13.8 39.3 53.9 33.7 17.5 37.6 32.1 18.1 9.4 23.0
@@ 4.2 Main Results
Table 1 compares the performance of ArrowGEV with state-of-the-art methods. As expected, models trained directly on the target benchmarks generally outperform zero-shot approaches. We further observe that RL-based methods consistently surpass those trained with SFT, likely because they optimize directly on temporal signals grounded in videos and the strong generalization ability of RL (Chu et al., 2025; Peng et al., 2025). In particular, ArrowGEV achieves higher accuracy than the strongest SFT-based baseline on all R1@m metrics. Moreover, compared to other RL-based methods, ArrowGEV yields average improvements in R1@m of 2.6% on Charades-STA, 1.9% on ActivityNet, and 2.5% on TVGBench. We attribute these gains to explicit training to understand the temporal structure, which enhances the robustness of the model in localizing events within videos.
@@ 4.3 OOD Generalization
Beyond GEV tasks, we further evaluate ArrowGEV on general video understanding and reasoning benchmarks. As shown in Figure 4, ArrowGEV yields consistent and significant improvements over the base Qwen2.5-VL-7B model across all benchmarks. These results highlight the versatility of our approach and demonstrate that explicitly modeling the arrow of time enhances OOD generalization across diverse video understanding and reasoning scenarios. We attribute this improvement to ArrowGEV’s ability to deeply understand time structure, which is a core element of general video understanding and reasoning.
@@
4.4 Improvement on Temporal directionality Understanding
As illustrated in Figure 5, ArrowGEV achieves a dramatic improvement over the base model in the time-sensitive and time-insensitive subsets. For time-sensitive subsets (upper row), the TDD score increases markedly, demonstrating that ArrowGEV successfully discriminates semantics that diverge under temporal reversal. Conversely, on the time-insensitive subset (bottom row), ArrowGEV achieves a significant reduction in TDD. This reduction confirms that the model maintains consistent localization across temporal flips, ensuring robust performance regardless of event directionality. These results indicate that ArrowGEV instills a deeper and more robust understanding of temporal directionality in VLMs. We further provide a case study in Appendix D.
@@ 4.5 Ablation Study
4.6 Performance on Different VLMs
To evaluate the versatility of our learning framework, we conducted experiments using different VLMs. Specifically, we examined the performance of our framework on Qwen-2.5-VL-Instruct-3B, and compared it with the best-performing RL baseline. Table 3 presents the mIoU metric on the benchmarks. The results indicate that ArrowGEV consistently improve the performance of the base model and outperform the best-performing baseline, indicating the effectiveness of our framework across different pre-trained VLMs. We report the performance on all metrics in Appendix C.
@@ 4.6 Performance on Different VLMs
Table 3: mIoU metric across three GE benchmarks.
@@ 5 Conclusion
We introduce ArrowGEV, which trains VLMs to enable both event grounding and video understanding through learning the temporal directionality. In addition to learning event localization in forward videos, ArrowGEV aims to distinguish the absence of time-sensitive events and recognize time-insensitive events in backward videos. Extensive experiments demonstrate that ArrowGEV outperforms various baselines on the event grounding task and improves general video understanding and reasoning performance compared to the base model.
@@
While our method demonstrates strong performance, the computational resources required in this model are expensive. Additionally, although our work primarily focuses on grounding events in videos, we believe that learning the Arrow of Time can be extended to more video reasoning tasks. We leave the exploration of this direction as future work.
@@
Appendix A Event Categorization Analysis
We use Qwen2.5-72B-Instruct (Yang et al., 2024) to categorize the type of events. To verify the reliability of the categorization, we watch the video and label 100 events from each dataset and compute the accuracy of the model prediction. The results in Table 5 indicate that LLMs’s prediction is highly aligned with humans, and can well recognize whether an event is time-sensitive or time-insensitive.
Appendix B Training Details
We leverage 7B and 3B models of Qwen2.5-VL (Bai et al., 2025) series as our base model. They are trained on large scale image and video data and show strong instruction following and reasoning abilities. During the post-training stage, we train the model for 5 epochs, and set a batch size of 128, learning rate 2e-5, number of candidate response $G=8$, and coefficient $\lambda=0.5$, KL term $\beta=0$, temperature in weight adjust $\tau=2$. We search these hyperparameters on the validation set. The checkpoint from the final epoch is used for all evaluations. All experiments were conducted on a single node equipped with 8 $\times$ H20 GPUs.
Appendix C Additional Results.
@@ Appendix B Training Details
Table 5: Event categorization performance on three benchmarks.
Appendix D Case Study
Figure 6: Case study on time-sensitive events (upper row) and time-insensitive events (bottom row). Although baseline methods can partially localize the correct timestamps in the forward video, they fail to distinguish the nonexistence of time-sensitive events in the backward videos and localize the time-insensitive events in the backward video.
@@ Appendix E Prompt
We list the prompt for event categorization in Figure 7.
Prompts for time sensitivity reasoning You are an AI assistant specializing in the analysis of temporal properties of events. You will be given a sentence describing an event. Your task is to: 1. Analyze the event described in the sentence. 2. Determine if the event is temporally sensitive or insensitive. 3. Output the results in a strict JSON format without any additional text or explanations. ### Input Event Sentence: \${sentence} ### Evaluation Criteria Time-Sensitive (sensitive: yes): The event has a clear forward direction. If played in reverse, it describes a different, often nonsensical or opposite, event. This indicates temporal asymmetry. Example: "A person puts a picture on the wall." (Reversed: "A person takes a picture off the wall.") Example: "A glass shatters." (Reversed: "Shards of glass assemble into a whole glass.") Time-Insensitive (sensitive: no): The event is a continuous state or a cyclical action. If played in reverse, the fundamental nature of the event does not change. This indicates temporal symmetry. Example: "A person is playing with a light switch." (Reversed: Still looks like a person playing with a light switch.) Example: "A ball is bouncing in place." (Reversed: Still a ball bouncing in place.) ### Output Format Now, please output your result below in a JSON format by filling in the placeholders in [] without any explanations: "reason": "[Briefly explain why the event is time-sensitive or time-insensitive, describing the forward and reverse action.]", "sensitive": "[yes/no]" """

### v2

@@ 4 Experiments
Benchmarks. We evaluate our model on three GEV benchmarks: Charades-STA (Sigurdsson et al., 2016), ActivityNet (Caba Heilbron et al., 2015), and TVGBench (Wang et al., 2025b). To further evaluate the generalization ability, we further compare ArrowGEV on the video understanding and reasoning benchmarks, including TempCompass (Liu et al., 2024), MVBench (Li et al., 2024b), VSI-Bench (Yang et al., 2025c), Video-MMMU (Hu et al., 2025), MMVU (Zhao et al., 2025), and VideoMME (Fu et al., 2024).
@@
where $\text{R1@m}(\text{fwd})$ denotes R1@m when predictions align with the ground truth $\mathcal{T}_{gt}$ on the forward video $\mathcal{V}$, and $\text{R1@m}(\text{rev})$ denotes R1@m on the backward video $\mathcal{V}^{\prime}$ with respect to the mirrored ground truth $\mathcal{R}(\mathcal{T}_{gt})$. The interpretation of the TDD depends on the event category. For Time-sensitive events, ideal models should accurately localize the forward event ($\text{R1@m}(\text{fwd})\rightarrow 1$) but recognize its absence upon reversal due to the nonexistence of the event ($\text{R1@m}(\text{rev})\rightarrow 0$), yielding a TDD approaching 1. For time-insensitive events, models should demonstrate temporal invariance by consistently localizing the event in both directions ($\text{R1@m}(\text{fwd})\approx\text{R1@m}(\text{rev})$), yielding a TDD approaching 0. This demonstrates that the model correctly recognizes the event’s invariance to temporal direction and exhibits strong consistency.
@@
Gemini-2.5-Flash (Comanici et al., 2025) 47.0 21.8 7.1 31.1 33.5 20.8 14.2 25.1 31.6 21.1 5.3 21.6
Gemini-2.5-Pro (Comanici et al., 2025) 53.9 25.5 8.8 34.6 48.2 31.9 18.7 36.3 40.0 25.7 11.4 29.1
GPT-5 (Singh et al., 2025) 39.7 18.3 6.2 28.4 48.2 33.0 18.8 35.9 31.2 18.8 6.3 22.3
@@
TVG-R1 (Chen et al., 2025b) 60.7 36.1 13.8 39.3 53.9 33.7 17.5 37.6 32.1 18.1 9.4 23.0
TimeLens-7B (Zhang et al., 2025) 70.7 39.8 14.5 42.3 53.5 35.2 19.7 37.7 38.3 21.7 12.2 25.7
GranAlign (Jeon et al., 2026) 59.1 39.6 22.7 38.0 50.3 34.0 16.5 33.1 - - - -
@@ 4.2 Main Results
Table 1 compares the performance of ArrowGEV with state-of-the-art methods. As expected, models trained directly on the target benchmarks generally outperform zero-shot approaches. We further observe that RL-based methods generally achieve competitive or superior performance compared with SFT-based ones, likely because they optimize directly on temporal signals grounded in videos and the strong generalization ability of RL (Chu et al., 2025; Peng et al., 2025). In particular, ArrowGEV achieves higher accuracy than the strongest SFT-based baseline on all R1@m metrics. Moreover, compared to other RL-based methods, ArrowGEV yields average improvements in R1@m of 2.6% on Charades-STA, 1.9% on ActivityNet, and 2.5% on TVGBench. We attribute these gains to explicit training to understand the temporal structure, which enhances the robustness of the model in localizing events within videos.
@@ 4.3 OOD Generalization
Beyond GEV tasks, we further evaluate ArrowGEV on general video understanding and reasoning benchmarks. As shown in Figure 4, ArrowGEV yields consistent and significant improvements over the base Qwen2.5-VL-7B model across all benchmarks. These results highlight the versatility of our approach and demonstrate that explicitly modeling the arrow of time enhances OOD generalization across diverse video understanding and reasoning scenarios. We attribute this improvement to ArrowGEV’s ability to deeply understand temporal structure, which is a core element of general video understanding and reasoning.
@@
4.4 Improvement on Temporal Directionality Understanding
As illustrated in Figure 5, ArrowGEV achieves a substantial improvement over the base model in the time-sensitive and time-insensitive subsets. For time-sensitive subsets (upper row), the TDD score increases markedly, demonstrating that ArrowGEV successfully discriminates semantics that diverge under temporal reversal. Conversely, on the time-insensitive subset (bottom row), ArrowGEV achieves a significant reduction in TDD. This reduction confirms that the model maintains consistent localization across temporal flips, ensuring robust performance regardless of event directionality. These results indicate that ArrowGEV instills a deeper and more robust understanding of temporal directionality in VLMs.
@@ 4.5 Ablation Study
Table 3: mIoU metric across three GEV benchmarks.
4.6 Performance on Different VLMs
To evaluate the versatility of our learning framework, we conducted experiments using different VLMs. Specifically, we examined the performance of our framework on Qwen-2.5-VL-Instruct-3B, and compared it with the best-performing RL baseline. Table 3 presents the mIoU metric on the benchmarks. The results indicate that ArrowGEV consistently improves the performance of the base model and outperforms the best-performing baseline, indicating the effectiveness of our framework across different pre-trained VLMs. We report the performance on all metrics in Appendix C.
Figure 6: Case study on time-sensitive events (upper row) and time-insensitive events (bottom row). Although baseline methods can partially localize the correct timestamps in the forward video, they fail to distinguish the absence of time-sensitive events in the backward video and fail to consistently localize time-insensitive events.
5 Case Study
@@ 6 Conclusion
We introduce ArrowGEV, which trains VLMs to enable both event grounding and video understanding through learning temporal directionality. In addition to learning event localization in forward videos, ArrowGEV aims to recognize the absence of time-sensitive events and recognize time-insensitive events in backward videos. Extensive experiments demonstrate that ArrowGEV outperforms various baselines on the event grounding task and improves general video understanding and reasoning performance compared to the base model.
@@
While our method demonstrates strong performance, the computational resources required by this model are substantial. Additionally, although our work primarily focuses on grounding events in videos, we believe that learning the Arrow of Time can be extended to more video reasoning tasks. We leave the exploration of this direction as future work.
@@
Appendix A Event Categorization Analysis
We use Qwen2.5-72B-Instruct (Yang et al., 2024) to categorize the type of events. To verify the reliability of the categorization, we watch the video and label 100 events from each dataset and compute the accuracy of the model prediction. The results in Table 5 indicate that LLM’s prediction is highly aligned with humans, and can well recognize whether an event is time-sensitive or time-insensitive.
@@ Appendix A Event Categorization Analysis
Table 5: Event categorization accuracy on three benchmarks.
Method Training time per step (s)
Time-R1 75.0
ArrowGEV 96.7
Table 6: Training time comparison.
Appendix B Training Details
We leverage the 7B and 3B models of Qwen2.5-VL (Bai et al., 2025) series as our base model. They are trained on large-scale image and video data and show strong instruction following and reasoning abilities. During the post-training stage, we train the model for 5 epochs, and set a batch size of 128, learning rate 2e-5, number of candidate responses $G=8$, and coefficient $\lambda=0.5$, KL term $\beta=0$, temperature in weight adjust $\tau=2$. We search these hyperparameters on the validation set. The checkpoint from the final epoch is used for all evaluations. All experiments were conducted on a single node equipped with 8 $\times$ H20 GPUs. See summary in Table 7.
Appendix C Additional Results.
Hyperparameter Value
Algorithm GRPO
Base Model Qwen2.5-VL-7B-Instruct
Max Prompt Length 8192
Max Response Length 1024
KL Coefficient($\beta$) 0
Reward Coefficient($\lambda$) 0.5
Learning Rate 2e-5
Sampling Temperature 1.0
Weighted Temperature $\tau$ 2.0
Batch Size 128
Rollout Number($G$) 8
Epochs 5
Table 7: Training Hyperparameters
Appendix D Training Efficiency
To compare training efficiencies, we recorded the average time required per training step for the Qwen2.5-VL-3B model. As detailed in Table 6, ArrowGEV incurs an approximate 28% increase in per-step training time due to the computational overhead of processing videos in both forward and backward directions.
@@ Appendix E Prompt
We list the prompt for event categorization and grounding in Figure 7 and Figure 8.
Prompt for time sensitivity categorization You are an AI assistant specializing in the analysis of temporal properties of events. You will be given a sentence describing an event. Your task is to: 1. Analyze the event described in the sentence. 2. Determine if the event is temporally sensitive or insensitive. 3. Output the results in a strict JSON format without any additional text or explanations. ### Input Event Sentence: \${sentence} ### Evaluation Criteria Time-Sensitive (sensitive: yes): The event has a clear forward direction. If played in reverse, it describes a different, often nonsensical or opposite, event. This indicates temporal asymmetry. Example: "A person puts a picture on the wall." (Reversed: "A person takes a picture off the wall.") Example: "A glass shatters." (Reversed: "Shards of glass assemble into a whole glass.") Time-Insensitive (sensitive: no): The event is a continuous state or a cyclical action. If played in reverse, the fundamental nature of the event does not change. This indicates temporal symmetry. Example: "A person is playing with a light switch." (Reversed: Still looks like a person playing with a light switch.) Example: "A ball is bouncing in place." (Reversed: Still a ball bouncing in place.) ### Output Format Now, please output your result below in a JSON format by filling in the placeholders in [] without any explanations: "reason": "[Briefly explain why the event is time-sensitive or time-insensitive, describing the forward and reverse action.]", "sensitive": "[yes/no]" """
@@
Prompt for event grounding To accurately pinpoint the event \${sentence} in the video, determine the precise time period of the event. Output your thought process within the <think> </think> tags. Then, provide the start and end times (in seconds, precise to two decimal places) in the format "start time to end time" within the <answer> </answer> tags. For example: "12.54 to 17.83".
Figure 8: The instruction to ground events in video.